OpenStudy (mitu12):

Solve: P= 2l+2w+2k for k

OpenStudy (anonymous):

First thing- is that 21, or 2L?

OpenStudy (mitu12):

L

OpenStudy (anonymous):

Alright, hold on...

OpenStudy (texaschic101):

P = 2L + 2w + 2k P - 2L - 2w = 2k (P - 2L - 2w) / 2 = k or P/2 - L - w = k

OpenStudy (mitu12):

so it will be k-L-w=k

OpenStudy (anonymous):

Well, there you go, lol. I was trying to draw it, but that's not the easiest thing in the world.

OpenStudy (mitu12):

ok thank you